# Non-Invasive Medical Digital Twins using Physics-Informed Self-Supervised Learning

### v1

@@ D.1 Deriving ODEs from circuit model
The (Simaan et al., 2008) model has five state variables which are determined by solving a system of ordinary differential equations (ODEs). We will derive five ODEs from basic circuit current laws. We obtain equations for the derivative of each of the five state variables. Recall
@@
The first row of the ODE system is given by conserving the current at the red node in Figure 6 as
@@
The second row is given as the following using the yellow circled node in Figure 6
@@
$x_{3}^{\prime}=(x_{2}-x_{3})/(R_{S}C_{S})+x_{5}/C_{S}.$
@@
$x_{4}^{\prime}=-x_{5}/C_{A}+\max(x_{1}-x_{4},0)/(R_{A}C_{A}).$
@@ D.2 Circuit Equivalence
We seek to find a reduced order equivalent circuit to improve model performance. We begin by ignoring diodes in our circuit to use linear equivalence, as then our circuit would have only linear components (capcitors, inductors, resistors), and we can use the follow result.
@@
For our case this gives us the following circuit, as we are interested in the capacitor representing pressure change in the left ventricle.
@@
We seek to write the variables of this new circuit in terms of the state variables in the (Simaan et al., 2008) model. To calculate the resistance of the equivalent circuits, we start by shorting all voltage sources (including capacitors) and opening all inductors. This gives us the following circuit.
@@
We want to leverage the reduced circuit architecture to solve for $P_{LV}(t)$. We know that the current through the load capacitor is $I_{cap}=(C(t)P_{LV}(t))^{\prime}$ using basic capacitor circuit relations. The Norton equivalent circuit gives a node to solve using Kirchoff’s law,
@@
Appendix E Validation with Mitral Stenosis labels on EchoNet
Figure 9: Overlayed PV loops for non-MS patients and MS patients with red patients being labeled as having MS and black patients being without labels. Abbreviations: MS=mitral stenosis.

### v2

@@ D.1 Deriving ODEs from circuit model
The Simaan et al. (65) model has five state variables which are determined by solving a system of ordinary differential equations (ODEs). We derive five ODEs from basic circuit current laws. We obtain equations for the derivative of each of the five state variables. Recall
@@
The first row of the ODE system is given by conserving the current $I$ at the red node in Figure 5 as
@@
The second row is given as the following by conserving current at the yellow circled node in Figure 5
@@
$x_{3}^{\prime}=(x_{2}-x_{3})/(R_{S}C_{S})+x_{5}/C_{S}$
by conserving current at the yellow node in Figure 5.
@@
$x_{4}^{\prime}=-x_{5}/C_{A}+\max(x_{1}-x_{4},0)/(R_{A}C_{A})$
by conserving current at the green node in Figure 5.
@@ D.2 Circuit Equivalence
We sought to find a reduced order equivalent circuit to improve model performance but were unsuccessful in finding an identifiable system. We began by ignoring diodes in our circuit to use linear equivalence, as then our circuit would have only linear components (capcitors, inductors, resistors), and we can use the follow result.
@@
For our case this gives us the circuit in Figure 6, as we are interested in the capacitor representing pressure change in the left ventricle.
@@
To identify this circuit, we wrote the variables of this new circuit in terms of the state variables in the (65) model. To calculate the resistance of the equivalent circuits, we start by shorting all voltage sources (including capacitors) and opening all inductors. This gives us the circuit in Figure 7.
@@
We would want to leverage the reduced circuit architecture to solve for $P_{LV}(t)$. We know that the current through the load capacitor is $I_{cap}=(C(t)P_{LV}(t))^{\prime}$ using basic capacitor circuit relations. The Norton equivalent circuit gives a node to solve using Kirchoff’s law,
@@
Appendix E Training details
The EchoNet model, achieving a Mean Absolute Error (MAE) of 5.40%, was trained for 3 hours over 13 epochs with a batch size of 100 and a learning rate of 0.001. In comparison, the CAMUS model, which reached an MAE of 6.81%, required 2 hours of training over 110 epochs with a batch size of 50 and a learning rate of 0.005.
In our experiments, we conducted a comprehensive hyperparameter search to ensure optimal performance. We employed a grid search approach to systematically explore a range of values for key hyperparameters. Specifically, we evaluated learning rates of 0.001, 0.005, and 0.0005, batch sizes of 50 and 100, and the number of epochs set to 300 and 500. We used the adam optimizer for training.
Models were trained on CPUs, each node featuring 24 cores (12 physical cores with hyper-threading) and 128 GB of RAM.
Appendix F Additional figures from experiments
F.1 Validation with Mitral Stenosis labels on EchoNet
Figure 8: Overlaid PV loops for non-MS patients and MS patients with red patients being labeled as having MS and black patients being without labels. Abbreviations: PV = pressure-volume, MS=mitral stenosis, LVV = Left-ventricular volume, LVP = Left-ventricular pressure.
F.2 Validation and comparison of physics-informed pretext task
Figure 9: Testing the pre-training task on a second synthetic dataset of 1000 randomly sampled points. The distribution of the loss is uniform across plausible parameter sets suggesting that our network is not subject to poor approximations in extreme cases.
Figure 10: Examples of patient-specific PINNs trained on synthetic data with 60 time points. Some PINNs learn the volume curve well. Others do not learn their patient curve after 100,000 epochs.
Figure 11: Examples of patient-specific Neural ODEs trained on synthetic data with 60 time points for each of the fives states in the cardiac model or trained only on volume state only.
Figure 12: Distribution of MAE on ejection fraction (EF) of the 20 best patient specific PINN and Neural ODE models used to create averages in Table 2.